# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v1

@@ 2 Basic tool
In this section, we give some useful tools
@@ 2.1 Littlewood-Paley theory
The first key tool is Littlewood-Paley decomposition. We introduce a operator named “nonhomogeneous dyadic blocks” $\Delta_{j}$ for $j\geq-1$, and satisfied for any $u$ belong to $\mathcal{S}^{\prime}(\mathbb{T}^{d})$ (Its also valid for $\mathbb{R}^{d}$), we have decomposition
@@
More precisely, its associated with two functions $\rho_{l}(k)$ and $\chi(k)$ under Fourier mode such that
@@
It can define $B_{p,r}^{s}(\mathbb{T}^{d})$ which contains Hölder space $\mathcal{C}^{\alpha}(\mathbb{T}^{d}):=B_{\infty,\infty}^{s}(\mathbb{T}^{d})$ for $\alpha\in(0,1)$ and Sobolev space $H^{s}(\mathbb{T}^{d}):=B_{2,2}(\mathbb{T}^{d})$ such that
@@
Define $f\prec g:=\sum_{j}S_{j-1}f\Delta_{j}g$, $f\succ g=g\prec f$, $f\circ g:=\sum_{|k-j|\leq 1}\Delta_{k}f\Delta_{j}g$, let $p,r\in[1,\infty]$ and $s\in\mathbb{R}$, then we have following estimates
@@
Let $s>0$, $\epsilon>0$, then following estimates hold:
@@
To explain the para-product, it must be notice that if $f$ and $g$ are two distribution, we can’t generally multiply them. A sufficient condition to do that is Lemma 2.1 holding which need $s_{1}+s_{2}\geq 0$ at least. Moreover, we call that $f\prec g$ and $f\succ g$ para-products since they are always well defined, and $f\circ g$ resonant term since this product limits the range in which their frequencies are close.
For our convenient, give following notation
@@
Note that $W^{\alpha}$ is a Banach algebra when $\alpha>0$.
@@
Sometimes it is also necessary to use a modified version of the paraproduct. Such operators have been introduced in [GIP15, GP17], and are defined as follows: Let $\varphi\in C^{\infty}(\mathbb{R},\mathbb{R}_{+})$ be nonnegative with compact support contained in $\mathbb{R}_{+}$ and with total mass 1, define $Q_{i}$ for all $i\geq 1$ as follows
$Q_{i}:C\mathcal{C}^{\beta}\to C\mathcal{C}^{\beta},\quad Q_{i}f(t)=\int_{\mathbb{R}}2^{-\gamma i}\varphi(2^{\gamma i}(t-s))f(s\wedge 0)\mathrm{d}s.$
With the help of $Q_{i}$, we define a modified para-product
@@
The remaining estimates will be presented in the Appendix B. Finally, we introduce a class of temporal regularity spaces, which have been referenced in [GIP15, GP17, Hai13, CC18]. Define the such norm with time $T$ as follows. $C_{T}X=C([0,T],X)$ for the space of continuous maps from $[0,T]$ to $X$, equipped with the norm $\sup_{t\in[0,T]}\|\cdot\|_{X}$. For $\alpha\in(0,1)$, we also define $C_{T}^{\alpha}X$ as a space of $\alpha$-Hölder continuous continuous functions from $[0,T]$ to $X$ equipped with the norm $\|f\|_{C_{T}^{\alpha}X}=\sup_{0\leq s<t\leq T}\frac{\|f_{s,t}\|_{X}}{|t-s|^{\alpha}}$. For our convenient, for $\alpha\geq 0$ we have following notation
@@
Its easy to check that $\mathcal{W}^{\alpha}(T)$ is a Banach algebra when $\alpha\geq 0$. For $\alpha<0$, we also give a definition $\mathcal{W}^{\alpha}_{p}(T)=C_{T}B_{p,p}^{\alpha}$ Next, we introduce some useful estimate
@@
Let $\alpha\in\mathbb{R}$, $\epsilon>0$, $\rho\in(0,\gamma)$ satisfied $\gamma\leq\rho+\epsilon$ and $\alpha+\rho<\gamma$. Let $f\in\mathcal{W}_{\infty}^{\epsilon}(T)$, $g\in C_{T}W^{\alpha+\rho}(T)$ and $\mathcal{L}g\in C_{T}W^{\alpha}$. Then
@@ 2.2 Stochastic tools
In this little section, we introduce some stochastic tools we need in this paper. The first tools is the Kolmogorov’s continuity criterion
@@
Then, there is a modification $\tilde{X}$ of $X$ whose sample paths are Hölder continuous with exponent $\alpha$ for every $\alpha\in(0,\frac{\epsilon}{p})$: This means that, for every $\omega\in\Omega$ and every $\alpha\in(0,\frac{\epsilon}{p})$, there exists a finite constant $C_{\alpha}(\omega)$ such that, for every $s,t\in I$,
@@
Now we give a basic analysis for white noise $\xi$. For the linear evolution $X(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s$, given a decomposition
$X(t)=\int_{-\infty}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s-e^{t\Lambda^{\gamma}}\int_{-\infty}^{0}e^{-s\Lambda^{\gamma}}\xi(s)\mathrm{d}s.$ (2.3)
Its not difficult to show that the second term is smooth, which means all the irregularity of $X(t)$ is described by the first term of the right hand side of (2.3). Hence we define $Y(t):=\int_{-\infty}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s$. For $X(t)$ and $Y(t)$, we have following lemma.
@@
The spatial Fourier transform $Y_{k}=\mathcal{F}_{x}Y(k)$ is a Gaussian process with zero-mean and satisfy
$\mathbb{E}[Y_{k}(t)Y_{k^{\prime}}(t^{\prime})]=e^{-|t-t^{\prime}||k|^{\gamma}}\frac{|\mathbb{T}|}{2|k|^{\gamma}}\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime}).$
where $|\mathbb{T}|=\int_{\mathbb{T}}1\mathrm{d}x$ and $\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime})$ is the indicator function. Moreover, for any $T>0$ and $\beta<\frac{\gamma(1-2\kappa)}{2}-\frac{1}{2}$, we have $X(t),Y(t)\in C_{T}^{\kappa}W^{\beta}$, for $\kappa<\frac{1}{2}$.
@@
Considering the second result, we can rewrite $Y(t)$ as
$Y(t)=\sum_{k\in\mathbb{Z}}Y_{k}(t)e^{ikx}.$
Calculate $\mathbb{E}|\Delta_{j}(Y(t)-Y(t^{\prime}))|^{2}$ and $\mathbb{E}\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2}}^{2p}$ as follows
$\mathbb{E}|\Delta_{j}(Y(t)-Y(t^{\prime}))|^{2}=$ $\mathbb{E}\sum_{k,k^{\prime}\in\mathbb{Z}}\rho_{j}(k)\rho_{j}(k^{\prime})[Y_{k}(t)-Y_{k}(t^{\prime})][Y_{-k^{\prime}}(t)-Y_{-k^{\prime}}(t^{\prime})]e^{i(k-k^{\prime})x}$
$=$ $\sum_{k,k^{\prime}\in\mathbb{Z}}\rho_{j}(k)\rho_{j}(k^{\prime})|\mathbb{T}|\frac{2-2e^{-|t-t^{\prime}||k|^{\gamma}}}{2|k|^{\gamma}}$
$\lesssim$ $\sum_{k\in\mathrm{supp}(\rho_{j})}k^{-\gamma(1-\delta)}|t-t^{\prime}|^{\delta}$
@@
for some $\delta\in[0,1]$. By Gaussian hypercontractivity [Jan97], we have
$\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2p}\lesssim(\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2})^{p}.$
@@
$\mathbb{E}\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2}}^{2p}=$ $\mathbb{E}(\int_{\mathbb{T}}|\Delta_{j}(Y(t)-Y(t^{\prime}))|^{2}\mathrm{d}x)^{p}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\mathbb{E}\prod_{i=1}^{p}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2p})^{\frac{1}{p}}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2})\mathrm{d}x_{i}$
@@
Then we can estimate $\mathbb{E}\|Y(t)-Y(t^{\prime})\|_{H^{\alpha}}^{2p}$ as
$\mathbb{E}\|Y(t)-Y(t^{\prime})\|_{B_{2,2}^{\beta}}^{2p}\lesssim$ $(\sum_{j\geq-1}2^{2j\beta}\mathbb{(}E\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2}}^{2p})^{\frac{1}{p}})^{p}$
$\leq$ $(\sum_{j\geq-1}2^{j(2\beta+1-\gamma(1-\delta))})^{p}|t-t^{\prime}|^{\delta p}<\infty,$
for $\beta<\frac{\gamma(1-\delta)-1}{2}$. By Lemma 2.6, we have $Y(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}H^{\frac{\gamma(1-\delta)-1}{2}}$. Choose $p$ large enough, we can prove the case of $H^{s}$. For Hölder space, we can also have the results by similar argument
$\mathbb{E}\|Y(t)-Y(t^{\prime})\|_{B_{2p,2p}^{\beta}}^{2p}=$ $\sum_{j\geq-1}2^{2j\beta p}\mathbb{E}\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2p}}^{2p}$
$\lesssim$ $\sum_{j\geq-1}2^{2j\beta p}\|\mathbb{E}|\Delta_{j}(Y(t)-Y(s))|^{2}\|_{L^{p}}^{p}$
$\lesssim$ $\sum_{j\geq-1}2^{jp(2\beta+1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta p}.$
Let $p$ large enough and by embedding for Besov space, we get $Y(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}\mathcal{C}^{\frac{\gamma(1-\delta)-1}{2}}$. Choosing $p$ large enough again, we can prove the case of $\mathcal{C}^{s}$, which we finish our proof. ∎
@@
By Lemma 2.7, we have $X(t),Y(t)\in C_{T}^{\kappa}W^{\beta}$ for $\frac{\alpha}{\gamma}<\frac{1}{2}-\frac{1}{\gamma}-\kappa$. Let $\beta=0$, then $\kappa$ can choose to be $\frac{1}{2}-\frac{1}{2\gamma}-\epsilon$ for any small $\epsilon>0$. Then we have $X(t),Y(t)\in\mathcal{W}^{\alpha}(T)$ a.s for any $T\geq 0$ and $\alpha<\frac{\gamma}{2}-\frac{1}{2}$ if we fix $\epsilon$ small enough.

### v2

@@ 2.1 Littlewood-Paley theory
The first key tool is Littlewood-Paley decomposition. We introduce an operator named “nonhomogeneous dyadic blocks” $\Delta_{j}$ for $j\geq-1$, and satisfying that for any $u$ belong to $\mathcal{S}^{\prime}(\mathbb{T}^{d})$ (It’s also valid for $\mathbb{R}^{d}$), we have decomposition
@@
More precisely, it’s associated with two functions $\rho_{l}(k)$ and $\chi(k)$ under Fourier mode such that
@@
It can define $B_{p,r}^{s}(\mathbb{T}^{d})$ which contains Hölder space $\mathcal{C}^{s}(\mathbb{T}^{d}):=B_{\infty,\infty}^{s}(\mathbb{T}^{d})$ for $\alpha\in(0,1)$ and Sobolev space $H^{s}(\mathbb{T}^{d}):=B^{s}_{2,2}(\mathbb{T}^{d})$ such that
@@
Define $f\prec g:=\sum_{j}S_{j-1}f\Delta_{j}g$, $f\succ g:=g\prec f$, $f\circ g:=\sum_{|k-j|\leq 1}\Delta_{k}f\Delta_{j}g$, let $p,r\in[1,\infty]$ and $s\in\mathbb{R}$, then we have the following estimates
@@
Let $s>0$, $\epsilon>0$, then the following estimates hold:
@@
To explain the paraproduct, we first note that two distributions $f$ and $g$ cannot, in general, be multiplied directly. A sufficient condition for such a product to be well defined is that the requirement of Lemma 2.1 be satisfied, which at least necessitates $s_{1}+s_{2}\geq 0$ . Moreover, $f\prec g$ and $f\succ g$ are called para-products since they are always well defined regardless of the regularity of $f$ and $g$. The term $f\circ g$ is referred to as the resonant term, because it involves interactions of frequencies that are close to each other, which typically restricts the range of regularities for which it can be controlled.
For our convenience, we give the following notation
@@
and note that $W^{\alpha}$ is a Banach algebra when $\alpha>0$.
@@
Sometimes it is also necessary to use a modified version of the paraproduct. Such operators have been introduced in [GIP15, GP17], and are defined as follows: Let $\varphi\in C_{c}^{\infty}(\mathbb{R},\mathbb{R}_{+})$ be nonnegative with compact support contained in $\mathbb{R}_{+}$ and with total mass 1, define $Q_{i}$ for all $i\geq 1$ as follows
$Q_{i}:C\mathcal{C}^{\beta}\to C\mathcal{C}^{\beta},\quad Q_{i}f(t)=\int_{\mathbb{R}}2^{\gamma i}\varphi(2^{\gamma i}(t-s))f(s\vee 0)\mathrm{d}s.$
We will often apply $Q_{i}$ and other operators on $C\mathcal{C}^{\beta}$ to functions $f\in C_{T}\mathcal{C}^{\beta}$ which we then simply extend from $[0,T]$ to R+ by considering $f(\cdot\wedge T)$. With the help of $Q_{i}$, we define a modified para-product
@@
Proof.
Noting that $\mathcal{F}(\mathcal{Q}_{j}S_{j-1}f(s)\Delta_{j}g(t))\in 2^{j}\tilde{\mathcal{C}}$, where $\tilde{\mathcal{C}}$ is some annuals, we are left with proving an appropriate estimate for $\|\mathcal{Q}_{j}S_{j-1}f(s)\Delta_{j}g(t)\|_{L^{p}}$. By Hölder inequality, and definition of $\varphi$, we have
$\|\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s\Delta_{j}g(t)\|_{L^{p}}\lesssim$ $\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))\|S_{j-1}f(s)\|_{L^{\infty}}\mathrm{d}s\|\Delta_{j}g(t)\|_{L^{p}}$
$\lesssim$ $\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}(t-s))\mathrm{d}s\|f\|_{C_{t}L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}}$
$\lesssim$ $\|f\|_{C_{t}L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}}.$
Since $g(t)\in W_{p}^{\beta}$, by Lemma 2.23 in [BCD11], we get the first estimate. The second estimate is similar, requiring only the observation that
$\sup_{j\geq-1}2^{j\alpha}\|S_{j-1}f(s)\|_{L^{\infty}}\leq C\|f(s)\|_{\mathcal{C}^{\alpha}},$ (2.2)
for $1<\alpha<0$ by Proposition 2.79 in [BCD11]. ∎
The remaining estimates will be presented in the Appendix B. Finally, we introduce a class of temporal regularity spaces, which have been referenced in [GIP15, GP17, Hai13, CC18]. Define the such norm with time $T$ as follows. $C_{T}X=C([0,T],X)$ for the space of continuous maps from $[0,T]$ to $X$, equipped with the norm $\sup_{t\in[0,T]}\|\cdot\|_{X}$. For $\alpha\in(0,1)$, we also define $C_{T}^{\alpha}X$ as a space of $\alpha$-Hölder continuous functions from $[0,T]$ to $X$ equipped with the norm $\|f\|_{C_{T}^{\alpha}X}=\sup_{0\leq s<t\leq T}\frac{\|f(t)-f(s)\|_{X}}{|t-s|^{\alpha}}$. For our convenience, for $\alpha\geq 0$ we have the following notation
@@
It’s easy to check that $\mathcal{W}^{\alpha}(T)$ is a Banach algebra when $\alpha\geq 0$. For $\alpha<0$, we also give a definition $\mathcal{W}^{\alpha}_{p}(T)=C_{T}B_{p,p}^{\alpha}$ Next, we introduce some useful estimate
@@
Let $\alpha\in\mathbb{R}$, $\epsilon,T>0$, $\rho\in(0,\gamma)$ satisfy $\gamma\leq\rho+\epsilon$ and $\alpha+\rho<\gamma$. Let $f\in\mathcal{W}_{\infty}^{\epsilon}(T)$, $g(0)=0$, $g\in C_{T}W^{\alpha+\rho}(T)$ and $\mathcal{L}g\in C_{T}W^{\alpha}$. Then
@@ 2.2 Stochastic tools
In this little section, we introduce some stochastic tools we need in this paper. The first tool is the Kolmogorov’s continuity criterion.
@@
Then, there is a modification $\tilde{X}$ of $X$ whose sample paths are Hölder continuous with exponent $\alpha$ for every $\alpha\in(0,\frac{\epsilon}{p})$. This means that, for every $\omega\in\Omega$ and every $\alpha\in(0,\frac{\epsilon}{p})$, there exists a finite constant $C_{\alpha}(\omega)$ such that, for every $s,t\in I$,
@@
Now we give a basic analysis for white noise $\xi$. For the linear evolution $X(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s$, we have the following lemma.
@@
The spatial Fourier transform $X_{k}=\mathcal{F}_{x}X(k)$ is a Gaussian process with zero-mean and satisfy
$\mathbb{E}[X_{k}(t)X_{k^{\prime}}(t^{\prime})]=(e^{-|t-t^{\prime}||k|^{\gamma}}-e^{-(t+t^{\prime})|k|^{\gamma}})\frac{|\mathbb{T}|}{2|k|^{\gamma}}\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime}),\quad\mathbb{E}[X_{0}(t)X_{k}(t^{\prime})]=(t\wedge t^{\prime})|\mathbb{T}|\textbf{I}_{k=0}(k).$
where $|\mathbb{T}|=\int_{\mathbb{T}}1\mathrm{d}x$ and $\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime})$ is the indicator function. Moreover, for any $T>0$ and $\beta<\frac{\gamma(1-2\kappa)}{2}-\frac{1}{2}$, we have $X(t)\in C_{T}^{\kappa}W^{\beta}$ for $\kappa<\frac{1}{2}$.
@@
Considering the second result, we can rewrite $X(t)$ as
$X(t)=\sum_{k\in\mathbb{Z}}X_{k}(t)e^{ikx}.$
For $j\geq 0(k\neq 0)$, we calculate $\mathbb{E}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}$ as follows
$\mathbb{E}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}=$ $\mathbb{E}\sum_{k,k^{\prime}\in\mathbb{Z}}\rho_{j}(k)\rho_{j}(k^{\prime})[X_{k}(t)-X_{k}(t^{\prime})][X_{-k^{\prime}}(t)-X_{-k^{\prime}}(t^{\prime})]e^{i(k-k^{\prime})x}$
@@
for some $\delta\in[0,1]$. The case j=1j=-1 can be treated using essentially the same arguments, except that then we need to distinguish the cases $k=0$ and $k\neq 0$. By Gaussian hypercontractivity [Jan97], we have
$\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2p}\lesssim(\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2})^{p}.$
@@
$\mathbb{E}\|\Delta_{j}(X(t)-X(t^{\prime}))\|_{L^{2}}^{2p}=$ $\mathbb{E}(\int_{\mathbb{T}}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}\mathrm{d}x)^{p}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\mathbb{E}\prod_{i=1}^{p}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2p})^{\frac{1}{p}}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2})\mathrm{d}x_{i}$
@@
Then we can estimate $\mathbb{E}\|X(t)-X(t^{\prime})\|_{H^{\beta}}^{2p}$ as
$\mathbb{E}\|X(t)-X(t^{\prime})\|_{B_{2,2}^{\beta}}^{2p}\lesssim$ $(\sum_{j\geq-1}2^{2j\beta}\mathbb{(}E\|\Delta_{j}(X(t)-X(t^{\prime}))\|_{L^{2}}^{2p})^{\frac{1}{p}})^{p}$
$\leq$ $(\sum_{j\geq-1}2^{j(2\beta+1-\gamma(1-\delta))})^{p}|t-t^{\prime}|^{\delta p}\leq C|t-s|^{\delta p},$
for $\beta<\frac{\gamma(1-\delta)-1}{2}$. By Lemma 2.6, we have $X(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}H^{\frac{\gamma(1-\delta)-1}{2}}$. Choose $p$ large enough, we can prove the case of $H^{s}$. For Hölder space, we can also have the results by similar argument
$\mathbb{E}\|X(t)-X(t^{\prime})\|_{B_{2p,2p}^{\beta}}^{2p}=$ $\sum_{j\geq-1}2^{2j\beta p}\mathbb{E}\|\Delta_{j}(X(t)-X(t^{\prime}))\|_{L^{2p}}^{2p}$
$\lesssim$ $\sum_{j\geq-1}2^{2j\beta p}\|\mathbb{E}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}\|_{L^{p}}^{p}$
$\lesssim$ $\sum_{j\geq-1}2^{jp(2\beta+1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta p}\leq C|t-s|^{\delta p}.$
Let $p$ be large enough and by embedding for Besov space, we get $X(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}\mathcal{C}^{\frac{\gamma(1-\delta)-1}{2}}$. Choosing $p$ large enough again, we can prove the case of $\mathcal{C}^{s}$, which we finish our proof. ∎
@@
By Lemma 2.7, we have $X(t)\in C_{T}^{\kappa}W^{\beta}$ for $\frac{\beta}{\gamma}<\frac{1}{2}-\frac{1}{\gamma}-\kappa$. Let $\beta=0$, then $\kappa$ can choose to be $\frac{1}{2}-\frac{1}{2\gamma}-\epsilon$ for any small $\epsilon>0$. Then we have $X(t)\in\mathcal{W}^{\alpha}(T)$ a.s for any $T\geq 0$ and $\alpha<\frac{\gamma}{2}-\frac{1}{2}$ if we fix $\epsilon$ small enough.